Determine if the following ratios form a proportion. Also, write the middle terms and extreme terms where the ratios form a proportion. 
25 cm : 1 m and ₹ 40 : ₹ 160 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

25 cm : 1 m and₹40 :₹160
25 cm = 25 / 100 m
= 0.25 m
0.25 / 1 = 1 / 4
40 / 160 = 1 / 4
Yes, these are in a proportion
Middle terms are 1 m,₹40 and Extreme terms are 25 cm,₹160

Thanks